Problem #2: A mining company is considering whether to develop a mining property. It is estimated that an immediate expenditure of S6 million will be needed to bring the property into production. Thereafter, the net cash inflow will be S1.4 million at the end of each year for the next 10 years. After that, an additional expenditure of S3 million at the end of the 11th year will be required to close down the mine and restore the surrounding area. For projects like this, the company would expect to earn at least j-2196. (a) Should the company proceed with the investment! (b) What would the immediate expenditure have to be before your answer to (a) would change? (A)No (B) Yes Problem #2(a): Select Problem #2(b): Cost in millions, correct to 2 decimals Just Save Submit Problem #2 for Grading Problem #2 | Attempt #1 Attempt #2 | Attempt #3 Attempt #4 2(a) 2(b) 2(a)2(a) Attempt #5 2(a) - 2(b) Your Answer: 2(a) 2(b) 2(a) 2(b) 2(a) 2(b) Your Mark:2(a) 2(b 2(a) 2(b) 2(a) 2(b)
Answer #1

Required Rate of Return = r = 21 %

Annual Cash Inflow = CF = $ 1.4 million

Immediate Initial Expenditure = IE = $ 6 million

Ending Closing Expense = ECE = $ 3 million

The project generates the aforementioned cash flows at the end of each year for a period of 10 years. Also, the closing expense is incurred at the end of year 11 and the immediate expense is incurred at end of year 0.

Therefore, Project NPV = - 6 + 1.4 x (1/0.21) x [1-{1/(1.21)^(10)}] - 3 / (1.21)^(11) = - $ 0.6928 or - $ 0.693 million

(a) As the project's NPV is negative it should not be taken up

(b) The immediate expenditure would have to be equal to the PV of the annual cash inflow less the PV of the ending closing expense, so that the project's NPV is zero.

Therefore, Required Immediate Expense = 1.4 x (1/0.21) x [1-{1/(1.21)^(10)}] - [3/(1.21)^(11)] = $ 5.30717 or approximately $ 5.31 million